Which of the following is the FORM of a particular solution of the differential equation y + 2y +y=tet Select one: O A. At
Answer #1

Given YA+qy+y=tet Yp tet Hitta 02720. tet (0+)2 = et t (D+1+1)2 = et 7 (0+2)2 7 - 11 et 2²(1+2)² (l+ 5)2+ et = 4 - et = = (1